OpenStudy (anonymous):

convert to a numeral in the base indicated 6 to a base 2

OpenStudy (anonymous):

6 =110 to the base 2

OpenStudy (anonymous):

do I get that by going like this 6/4 = 1 with remainder 2 2/2 = 1 with remainder 0 110 base 2

OpenStudy (anonymous):

correct method is to divide 6 by 2 repeatedly and get the remainder

OpenStudy (anonymous):

then when the remainder is 0 do I always put that on the end

OpenStudy (anonymous):

of my answer

OpenStudy (anonymous):

do it like this 6/2 = 3 with remainder = 0 3/2= 1 with remainder =1 so the ans is 110

OpenStudy (anonymous):

so when do you know when to stop dividing I am thinking when the answer is smaller then the base

OpenStudy (anonymous):